Dennis Murphy is a salaried, nonexempt administrative assistant for Dionti Investments and is paid semimonthly. He is married with two withholding allowances. His annual salary is $70,200, and his standard workweek is 37.5 hours. During the pay period, he worked 10 hours overtime.

Required:
Compute Dennis's gross pay for the period ending August 18, 20XX. (Do not round intermediate computations. Round your final answers to 2 decimal places.)

Company Dionti Investments Period Ended: Aug 18, 20XX
Name M/S # W/H Period Wage Hourly Rate No. of Overtime Hours Regular Earnings Overtime Earnings Gross Earnings
Dennis Murphy
Answer #1

Answer :

Company Dionti Investments Period Ended: Aug 18, 20XX
Name M/S # W/H Period Wage Hourly Rate No. of Overtime Hours Regular Earnings Overtime Earnings Gross Earnings
Dennis Murphy married 36 10 $2,825 $ 540 $3,465

1. Calculation of overtime earnings :

Annual salary = $70,200

No. Of weeks = 52 weeks

Weekly salary = (70200/52) = $1,350

Weekly hrs = 37.5 hrs

Hourly rate = (1,350 / 37.5) = $ 36 per hr

Overtime rate = 36 × 1.5 = $ 54 per hr

Total overtime earnings = 54 × 10 = $540

2. Calculation of regular earnings

Regular earnings paid semimonthly =( 70200/12) ÷ 2 = $2,925

3. Calculation of gross pay

Gross pay for period ending Aug 18, 20xx =

Regular Earnings + Overtime Earnings

= 540 + 2925 = $ 3,465